# Revealing Language Model Trajectories via Kullback-Leibler Divergence

### v1

@@ Appendix B Related Work and Discussions on Diffusion Exponents
In Section 3, we estimated the diffusion exponents $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ corresponding to the trajectories of models in weight and log-likelihood spaces, respectively. Our analysis revealed that during the early phase of training, both exponents are relatively large ($c_{\mathrm{w}}=1.2$ and $c_{\mathrm{q}}=0.8$), but they decrease as training progresses. At the later stages, the exponent in weight space approaches $c_{\mathrm{w}}\approx 1$, which corresponds to Brownian motion, while the exponent in log-likelihood space becomes significantly smaller ($c_{\mathrm{q}}\ll 1$), which is characteristic of subdiffusion.
@@
In our experimental setting, texts were sampled from a corpus that closely approximates the training data. Consequently, each element of the log-likelihood vector represents a term contributing to the cross-entropy loss, and the negative average of these terms serves as an unbiased estimator of the expected loss. The contrast between the relatively large diffusion exponent in weight space and the much smaller exponent in log-likelihood space suggests that even as model weights continue to change, the output probabilities remain stable. As a result, the loss function is largely unaffected by these weight changes. This observation aligns with the hypothesis that stochastic gradient descent favors flat minima Keskar et al. (2017).
Moreover, near the loss minima, the sum of negative log-likelihoods is constrained as cross-entropy loss, implying convergence to regions where many texts have high log-likelihood values. The observed small diffusion exponent suggests that the movement of the vectors becomes restricted in these regions.
Appendix C Standard Error of the Estimated KL Divergence
@@
Let $\bm{q}_{i}$ and $\bm{q}_{j}$ denote the $\bm{q}$ coordinates of models $p_{i}$ and $p_{j}$, respectively. While we use $\bm{q}$ pre-scaled by $1/\sqrt{N}$ in the main body of this paper, we do not apply this scaling in this appendix. Accordingly, the KL divergence between the models is estimated as $\mathrm{KL}(p_{i},p_{j})\approx\frac{1}{N}\|\bm{q}_{i}-\bm{q}_{j}\|^{2}=\frac{1}{N}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{2},$ which corresponds to the sample mean of $\{(Q_{ik}-Q_{jk})^{2}\}_{k=1}^{N}$. Assuming that the effects of double-centering in $\bm{Q}$ are negligible, these terms can be treated as approximately independent. When $N$ is sufficiently large, the variance of the estimated KL divergence is given by
@@
$\approx\frac{1}{N}\left\{\frac{1}{N}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{4}-\frac{1}{N^{2}}\|\bm{q}_{i}-\bm{q}_{j}\|^{4}\right\}$
$\approx\frac{1}{N^{2}}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{4}-\frac{1}{N}\mathrm{KL}(p_{i},p_{j})^{2}.$
@@
$\approx\sqrt{\frac{1}{N^{2}}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{4}-\frac{1}{N}\mathrm{KL}(p_{i},p_{j})^{2}}.$
Appendix D Details of Section 3
@@
To analyze the trajectories of language models during pretraining, we used Pythia models [^6] Released under the Apache 2.0 License. of various sizes, as listed in Table LABEL:tab:model_list_pretraining in Appendix G. For each checkpoint, we computed log-likelihood vectors using the same set of 10,000 texts as in Oyama et al. (2025). The computations were performed on an NVIDIA RTX 6000 Ada and took approximately 10 minutes for a 7B model loaded in float16 precision.
D.1 Preprocessing
@@
Figure 8: Squared Euclidean distance between the weights of consecutive checkpoints for Pythia 1B. The distances between 115k and 116k, and between 116k and 117k are abnormally large.
@@
Figure 9: The relationship between the maximum and the standard deviation of the absolute differences in log-likelihoods between consecutive checkpoints of the same model size, computed for each text.
@@
Figure 10: Change in KL divergence after step 10k as a function of the number of top outlier texts removed (10, 100, 200, …, 1000). Brighter colors indicate larger numbers of removed texts.
@@
Assuming that the KL divergence between consecutive checkpoints changes smoothly, we regarded certain checkpoints and texts as outliers. All experiments were conducted after removing these outliers. This appendix provides a detailed description of the preprocessing procedure.
@@
Following Oyama et al. (2025), we clip the log-likelihood matrix $\bm{L}$ at the bottom 2%, after excluding the outlier model. Next, we consider the log-likelihood difference for each text between consecutive checkpoints after the warmup phase. For each text, we define its outlier score as the maximum absolute difference across all model sizes and training steps. Specifically, letting $p_{i,t}$ denote the model of size $i$ at step $t$, the outlier score for a text $x$ is given by $\max_{i,t}\left\{\left|p_{i,t+1}(x)-p_{i,t}(x)\right|\right\}$. It is also possible to assign scores to texts using the standard deviation of these differences. However, Fig. 9 shows that the maximum and the standard deviation are highly correlated. Therefore, we adopt the maximum value for scoring. The figure further confirms the presence of texts that clearly behave as outliers. For example, the text with the highest outlier score was a meaningless and repetitive sequence of symbols, such as 028a28a0028a28a0028a28a....
Fig. 10 shows the KL divergence between consecutive checkpoints after step 10k [^8] The maximum is computed from step 2k onward, after the warmup phase, but we visualize from step 10k to better capture the variation in KL divergence., plotted as a function of the number of top outlier-scored texts removed (top 10, 100, 200, …, 1000). When the top 300 texts (3%) are removed, the variation in KL divergence stabilizes with respect to the number of removed texts. Therefore, we excluded these 300 texts from our experiments.
D.2 KL Divergence between Checkpoints
(a) KL divergence between consecutive checkpoints during the warmup phase. The horizontal axis is scaled as $\log_{10}(1+\mathrm{Step})$.
(b) KL divergence between consecutive checkpoints after the warmup phase. When the vertical axis is plotted on a logarithmic scale, this corresponds to the right panel of Fig. 4.
Figure 11: KL divergence between consecutive checkpoints of Pythia. For example, the value at step 10k on the horizontal axis represents the KL divergence between the checkpoints at steps 10k and 11k.
Figure 12: KL divergence and its standard error between consecutive checkpoints of Pythia after step 10k.

### v2

@@ Appendix B Related Work and Discussions on Diffusion Exponents
Anomalous diffusion in weight space.
It has been reported that, during the training of neural networks, trajectories in weight space can exhibit anomalous diffusion, and this phenomenon has recently attracted considerable research interest Chen et al. (2022); Kunin et al. (2024); Ly and Gong (2025). These studies analyze training dynamics through the temporal scaling of distances in weight space and report deviations from standard Brownian motion.
Diffusion exponents in weight and log-likelihood spaces.
@@
In Section 4, we estimated the diffusion exponents $c_{\mathrm{w}}$ and $c_{\mathrm{q}}$ for trajectories in weight space and log-likelihood space, respectively. Our results show that during the early phase of training, both exponents take relatively large values ($c_{\mathrm{w}}=1.2$ and $c_{\mathrm{q}}=0.8$), but decrease as training progresses. In the later stages, the exponent in weight space approaches $c_{\mathrm{w}}\approx 1$, corresponding to Brownian-like diffusion, whereas the exponent in log-likelihood space becomes much smaller ($c_{\mathrm{q}}\approx 0.2\ll 1$), which is characteristic of subdiffusion. In our discussion, fractional Brownian motion is used purely as an interpretive model to relate the observed diffusion exponents to effective fractal dimensions, rather than as a generative model of the training dynamics.
We note that the estimated diffusion exponent $c_{\mathrm{q}}$ exhibits systematic variation across model sizes, as shown in the right panel of Fig. 5. In particular, smaller models tend to show lower values ($c_{\mathrm{q}}\approx 0.1\text{--}0.2$), while larger models exhibit higher values ($c_{\mathrm{q}}\approx 0.2\text{--}0.3$). While we report $c_{\mathrm{q}}\approx 0.2$ as a representative value capturing the typical subdiffusive behavior in log-likelihood space, a more detailed analysis of model-size dependence is beyond the scope of this paper.
Interpretation via loss landscape and flat minima.
@@
In our experimental setting, texts are sampled from a corpus that closely approximates the training data. Consequently, each component of the log-likelihood vector corresponds to a term contributing to the cross-entropy loss, and the negative average of these components serves as a reasonable empirical estimate of the expected loss. The marked contrast between the near-Brownian diffusion in weight space and the strongly suppressed diffusion in log-likelihood space indicates that, even as model weights continue to change, the output probability distributions remain relatively stable. This observation is consistent with the hypothesis that stochastic gradient descent tends to favor flat regions of the loss landscape Keskar et al. (2017).
Geometric interpretation via folding.
In addition to such optimization-based explanations, a more structural interpretation arises from the intrinsic redundancies of multilayer neural networks. Due to permutation symmetries of hidden units, many distinct weight configurations correspond to the same function or output distribution Fukumizu and Amari (2000). From a statistical viewpoint, such permutation-induced redundancies are associated with degenerate directions of the Fisher information matrix, giving rise to flat directions in the loss landscape. As a result, trajectories that are well separated in weight space can be mapped to nearby or identical points in the log-likelihood space. Viewed through this many-to-one mapping, training trajectories in the log-likelihood space are strongly folded, leading to complex geometric structures that remain confined to a relatively narrow region of the space.
Quantifying folding via Hölder regularity.
The geometric interpretation based on folding can be further quantified by considering the regularity of the mapping $f:\bm{W}\to\bm{q}$ from weight space to log-likelihood space. Specifically, recall that a mapping is $\alpha$-Hölder continuous if distances in the output space are bounded by a constant times the $\alpha$-th power of distances in the input space, with $\alpha=1$ corresponding to Lipschitz continuity and smaller values of $\alpha>0$ indicating increasing non-smoothness. If $f$ is $\alpha$-Hölder continuous and satisfies the non-degeneracy condition required to saturate the theoretical bound, then the fractal dimensions of the corresponding trajectories obey $D_{\mathrm{q}}=D_{\mathrm{w}}/\alpha$ Falconer (2014), which implies
$\alpha=\frac{D_{\mathrm{w}}}{D_{\mathrm{q}}}=\frac{c_{\mathrm{q}}}{c_{\mathrm{w}}}.$
This relation shows that as $\alpha$ decreases, the mapping can strongly fold the weight-space trajectory into a geometrically more complex and higher-dimensional trajectory in the log-likelihood space, while its large-scale displacement remains strongly suppressed. In our empirical setting, the observed diffusion exponents imply effective dimensions $D_{\mathrm{w}}\approx 2$ and $D_{\mathrm{q}}\approx 10$, which in turn yield an effective Hölder exponent of $\alpha\approx 0.2$.
A constructive analogue for folding in log-likelihood space.
Although the mapping $f$ induced by a neural network is explicitly defined by its architecture and parameters, its geometric regularity and folding structure are not directly accessible in closed form. This folding mechanism can nevertheless be illustrated by a constructive example provided by generalized Takagi–Landsberg functions. These functions are defined as multiscale summations of sawtooth functions $S(x)=\text{dist}(x,\mathbb{Z})$:
$f(\bm{W})=\sum_{k=1}^{\infty}\mathbf{a}_{k}\lambda^{-k\alpha}S\!\left(\lambda^{k}\mathbf{b}_{k}^{\top}\bm{W}\right),$
where $\lambda>1$ is a constant and $\mathbf{a}_{k}$ and $\mathbf{b}_{k}$ are vectors with the same dimensions as $\bm{q}$ and $\bm{W}$, respectively. This class of functions is known to be $\alpha$-Hölder continuous for any $0<\alpha<1$ under mild boundedness and non-degeneracy conditions on $\{\mathbf{a}_{k},\mathbf{b}_{k}\}$ Shidfar and Sabetfakhri (1986); Allaart and Kawamura (2011). Such constructions provide a simple geometric illustration of how the many-to-one and redundant parameterization of deep neural networks can give rise to highly folded trajectories in the log-likelihood space. In contrast to this construction, previous studies have focused on folding phenomena in input–output mappings of deep networks, where depth induces increasingly fine sawtooth partitions of the input space Telgarsky (2016, 2017).
Appendix C On the Estimation of the KL Divergence
Interpretation of KL divergence via text entropy.
As a reference for interpreting KL divergence values, we estimate the entropy of the text from log-likelihood statistics, as described below. In our setting, this estimated text entropy is approximately 0.71 bits/byte; for example, a KL divergence of 0.1 bits/byte corresponds to a value slightly over 10% of the text entropy.
In this work, we define the minimum value of the negative mean log-likelihood across models as an estimate of the text entropy. This is because, given a model $p_{i}$, a text set $\{x_{1},\ldots,x_{N}\}$, the log-likelihood $\ell_{i}(x_{s})$, and the true data-generating distribution $p_{0}$, the following approximation holds:
$-\frac{1}{N}\sum_{s=1}^{N}\ell_{i}(x_{s})$
$\quad\approx\mathbb{E}_{x\sim p_{0}}[-\log p_{i}(x)]$
$\quad=\mathbb{E}_{x\sim p_{0}}[-\log p_{0}(x)]+\mathrm{KL}(p_{0},p_{i}).$
Therefore, the minimum of the negative mean log-likelihood across models can be considered a good approximation of the text entropy.
Standard error of the estimated KL divergence.
@@ Appendix C On the Estimation of the KL Divergence
Let $\bm{q}_{i}$ and $\bm{q}_{j}$ denote the $\bm{q}$ coordinates of models $p_{i}$ and $p_{j}$, respectively. While in the main text of this paper we use $\bm{q}$ coordinates rescaled by $1/\sqrt{2N\bar{B}\log 2}$ so that squared Euclidean distances approximate KL divergence in bits per byte, we do not apply this rescaling in this appendix. Accordingly, the KL divergence between the models is estimated as $\mathrm{KL}(p_{i},p_{j})\approx\frac{1}{2N}\|\bm{q}_{i}-\bm{q}_{j}\|^{2}=\frac{1}{2N}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{2},$ which corresponds to the sample mean of $\{(Q_{ik}-Q_{jk})^{2}/2\}_{k=1}^{N}$. Assuming that the effects of double-centering in $\bm{Q}$ are negligible, these terms can be treated as approximately independent. When $N$ is sufficiently large, the variance of the estimated KL divergence is given by
@@
$\approx\frac{1}{N}\left\{\frac{1}{4N}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{4}\right$.
$\qquad\left.-\frac{1}{4N^{2}}\|\bm{q}_{i}-\bm{q}_{j}\|^{4}\right\}$
$\approx\frac{1}{4N^{2}}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{4}-\frac{1}{N}\mathrm{KL}(p_{i},p_{j})^{2}.$
@@
$\approx\sqrt{\frac{1}{4N^{2}}\sum_{k=1}^{N}(Q_{ik}-Q_{jk})^{4}-\frac{1}{N}\mathrm{KL}(p_{i},p_{j})^{2}}.$
The standard error estimated in nats above is converted to bits per byte by dividing by $\bar{B}\log 2$. Equivalently, when using $\bm{q}$ coordinates rescaled by $1/\sqrt{2N\bar{B}\log 2}$ as in the main text, the standard error of KL divergence in bits per byte is given by
$\mathrm{SE}(\widetilde{\mathrm{KL}}(p_{i},p_{j}))$
$\approx\sqrt{\sum_{k=1}^{N}(\tilde{Q}_{ik}-\tilde{Q}_{jk})^{4}-\frac{1}{N}\widetilde{\mathrm{KL}}(p_{i},p_{j})^{2}},$
where $\tilde{Q}=Q/\,\sqrt{2N\bar{B}\log 2}$ and
$\widetilde{\mathrm{KL}}(p_{i},p_{j})=\frac{\mathrm{KL}(p_{i},p_{j})}{\bar{B}\log 2}=\sum_{k=1}^{N}(\tilde{Q}_{ik}-\tilde{Q}_{jk})^{2}.$
On the robustness to the sampling of texts.
As an additional experiment, we computed the log-likelihood vectors using a new set of 10,000 texts randomly sampled from the 5,703,791 texts in a subset of the Pile [^6] The full Pile corpus, when segmented into 1,024-byte text chunks, yields a total of 5,703,791 texts Oyama et al. (2025)., using the Pythia-410M model. We then computed the pairwise KL divergence between all checkpoints after warmup, and found that the Pearson correlation coefficient between the KL values obtained using the original text set and those obtained using the new text set was 0.99. This indicates that the impact of text sampling on the KL values is limited.
One advantage of using log-likelihood vectors is that the resolution for each domain can be adjusted by the amount of text from that domain included in the dataset. If a specific domain is fixed, it effectively corresponds to projecting the trajectory onto a certain aspect, and thus the shape of the trajectory is expected to change.
Appendix D Additional Visual Analyses of Pretraining Trajectories
This appendix provides supplementary visual analyses of pretraining trajectories. Appendix D.1 and Appendix D.2 focus on trajectories in the log-likelihood space, using three-dimensional t-SNE and PCA to examine their continuity and global structure, respectively. In contrast, Appendix D.3 presents trajectory visualizations in the weight space, highlighting fundamental limitations of joint visualization across models.
D.1 Three-Dimensional t-SNE Visualizations
(a) Visualization of pretraining trajectories for five Pythia model sizes using 3D t-SNE (perplexity = 30). This figure corresponds to a three-dimensional version of Fig. 1, excluding additional random seeds of the 410M model.
(b) PC2 versus PC3 for the later portion of the Pythia-2.8B training trajectory after applying 3D t-SNE.
Figure 6: Pythia training trajectories visualized using 3D t-SNE.
Figure 7: Autocorrelation functions of PC3 for the later portion of the Pythia-2.8B training trajectory (from step 67k onward) after applying 3D t-SNE, shown for different perplexity values.
This subsection presents three-dimensional t-SNE visualizations of the pretraining trajectories discussed in Section 4. In the main text, Fig. 1 presents two-dimensional t-SNE visualizations of pretraining trajectories for Pythia models across multiple model sizes and random seeds. When multiple trajectories are embedded into two dimensions, apparent discontinuities may arise near regions where trajectories overlap. This behavior reflects a geometric limitation of low-dimensional embeddings: in general, embedding multiple curves without crossings requires a space of dimension three or higher.
As shown in Fig. 6, this issue is resolved by visualizing the trajectories using three-dimensional t-SNE. In the 3D embedding, each training trajectory can be visualized as a continuous curve without crossings, supporting the continuity of trajectories in the original high-dimensional log-likelihood space. For clarity, we exclude additional seeds of the Pythia-410M model and visualize the trajectories of the five model sizes simultaneously in three dimensions, as shown in Fig. 6.
Figure 6 further visualizes the later portion of the Pythia-2.8B trajectory by plotting PC2 versus PC3 after applying 3D t-SNE. In this representation, the trajectory appears to exhibit a spiral-like structure. To analyze this phenomenon, Fig. 7 shows the autocorrelation function of PC3 for the later portion of the trajectory, computed under different t-SNE perplexity settings (20, 30, 40, and 50). The spiral period is estimated as the training step at which the autocorrelation function attains its maximum between the second and third zero-crossings. As the perplexity increases, the estimated period also increases, indicating that the apparent spiral structure depends on the neighborhood size used in the embedding.
This spiral-like appearance can be attributed to the geometric constraint of embedding high-dimensional, intricately folded trajectories into three dimensions without crossings. Such folding is consistent with the fractal-like structure suggested by the power-law scaling of the training trajectories discussed in the main text. Accordingly, the spiral pattern should be interpreted as a visualization artifact rather than as an intrinsic property of the training dynamics.
D.2 PCA-Based Visualizations of Pretraining Trajectories
We further analyze the same pretraining trajectories using principal component analysis (PCA) to highlight their global geometric structure. In contrast to t-SNE, which emphasizes local neighborhood relationships, PCA captures large-scale variations shared across different model sizes.
Fig. 8 shows the result of applying PCA simultaneously to all model sizes and plotting their trajectories. Fig. 9 further illustrates the evolution of multiple principal component values along the training steps. Unlike t-SNE, the top principal components exhibit similar values across model sizes, resulting in overlapping trajectories. This difference likely arises from the fact that the top principal components capture global variation in the trajectories, whereas t-SNE emphasizes local structure. In contrast, the lower principal components exhibit more noise-like behavior. As shown in Table 2, the KL divergence between different model sizes at the final checkpoint is larger than that between consecutive saved checkpoints. This is consistent with the observation that trajectories are separated in the lower components.
D.3 Limitations of Trajectory Visualization in Weight Space
Figure 10: t-SNE visualization (perplexity=30) of the trajectories based on pairwise distances in the weight space for Pythia models after step 100k. To align plot ranges across model sizes, the trajectories were centered after t-SNE. They were also rotated to roughly align the directions of progression. Final checkpoints are indicated using distinct markers.
Since pretraining trajectories in the weight space do not admit a joint visualization across different model sizes and random seeds, we visualize them separately for each model. For each model size, we computed a distance matrix containing the pairwise Euclidean distances between the weight vectors of all checkpoints. Figure 10 shows the trajectories from step 100k onward, visualized using t-SNE777t-SNE can take a distance matrix as input.. The trajectories were centered after t-SNE to align the plot ranges. In addition, for the 1B model, the trajectory was rotated by 180 degrees to roughly align its direction of progression with those of the other model sizes.
Despite these adjustments, the resulting visualizations remain difficult to interpret. Because each model requires a separate embedding, the relative positions and orientations of trajectories across models are arbitrary and cannot be meaningfully compared. This highlights a fundamental limitation of weight-space representations for trajectory visualization, in contrast to the log-likelihood space, which provides a shared coordinate system for joint analysis.
Appendix E Details of Sections 3.1 and 4
@@ D.3 Limitations of Trajectory Visualization in Weight Space
To analyze the trajectories of language models during pretraining, we used Pythia models [^8] Released under the Apache 2.0 License. of various sizes, as listed in Table LABEL:tab:model_list_pretraining in Appendix I. For each checkpoint, we computed log-likelihood vectors using the same set of 10,000 texts as in Oyama et al. (2025). The computations were performed on an NVIDIA RTX 6000 Ada and took approximately 10 minutes for a 7B model loaded in float16 precision.
E.1 Preprocessing
@@
Figure 11: Squared Euclidean distance between the weights of consecutive saved checkpoints for Pythia 1B. The distances between 115k and 116k, and between 116k and 117k are abnormally large.
@@
Figure 12: The relationship between the maximum and the standard deviation of the absolute differences in log-likelihoods between consecutive saved checkpoints of the same model size, computed for each text.
@@
Figure 13: Change in KL divergence after step 10k for different numbers of top outlier texts removed (10, 100, 200, …, 1000). Brighter colors indicate larger numbers of removed texts.
@@
Assuming that the KL divergence between consecutive saved checkpoints changes continuously, we regarded certain checkpoints and texts as outliers. All experiments were conducted after removing these outliers. This appendix provides a detailed description of the preprocessing procedure.
@@
Following Oyama et al. (2025), we clip the log-likelihood matrix $\bm{L}$ at the bottom 2%, after excluding the outlier model. Next, we consider the log-likelihood difference for each text between consecutive saved checkpoints after the warmup phase. For each text, we define its outlier score as the maximum absolute difference across all model sizes and training steps. Specifically, letting $p_{i,t}$ denote the model of size $i$ at step $t$, the outlier score for a text $x$ is given by $\max_{i,t}\left\{\left|\log p_{i,t+1}(x)-\log p_{i,t}(x)\right|\right\}$. It is also possible to assign scores to texts using the standard deviation of these differences. However, Fig. 12 shows that the maximum and the standard deviation are highly correlated. Therefore, we adopt the maximum value for scoring. The figure further confirms the presence of texts that clearly behave as outliers. For example, the text with the highest outlier score was a meaningless and repetitive sequence of symbols, such as 028a28a0028a28a0028a28a....
Fig. 13 shows the KL divergence between consecutive saved checkpoints after step 10k [^10] The maximum is computed from step 2k onward, after the warmup phase, but we visualize from step 10k to better capture the variation in KL divergence., shown for different numbers of top outlier-scored texts removed (top 10, 100, 200, …, 1000). When the top 300 texts (3%) are removed, the variation in KL divergence stabilizes with respect to the number of removed texts. Therefore, we excluded these 300 texts from our experiments.
E.2 KL Divergence between Checkpoints
Figure 14: KL divergence and its standard error between consecutive saved checkpoints of Pythia after step 10k.